OpenStudy (anonymous):

WILL FAN AND MEDAL ONE EASY GEOMETRY QUESTION! Which of the following is the equation of a circle with a radius of 2.5 and its center at (4, –3)? (x – 4)2 + (y + 3)2 = 2.5 (x + 4)2 + (y – 3)2 = 2.5 (x – 4)2 + (y + 3)2 = 6.25 (x + 4)2 + (y – 3)2 = 6.25

OpenStudy (twizttiez):

What do you think?

OpenStudy (anonymous):

@TwiztTiez , I think its option 2?

OpenStudy (twizttiez):

I don't think so

OpenStudy (anonymous):

@TwiztTiez I dont know how to solve it, im stuck:( my second guess would be option one

OpenStudy (twizttiez):

Remember the general circle equation: (x - h)² + (y - k)² = r² Where: h = x coordinate of the vertex k = y coordinate of the vertex r = radius Given: radius of 2.5 Means: r = 2.5 Given: center at (-4, 3) Means: h = -4 Means: k = 3 Plug these into the general equation. (x - h)² + (y - k)² = r² (x - -4)² + (y - 3)² = (2.5)² (x + 4)² + (y - 3)² = 6.25 ANSWER: D. (x + 4)² + (y - 3)² = 6.25 Via~ Yahoo Answers

OpenStudy (twizttiez):

Now please study over this so you'll understand what to do

OpenStudy (anonymous):

Thank you so much, I will! :) @TwiztTiez

OpenStudy (twizttiez):

No prob